Can somebody answer this quiz ?

There are four cities formed a square.

We have to build a road. On the road we have to reach from any city to any other city.

We have money for only 22 km road, exactly . So, how to build ??
1
Expert's answer
2020-04-17T18:39:31-0400

Four cities formed a square.We have to build a road. On the road we have to reach from any city to any other city.

We have money for only 22 km road, exactly .

"4a+2\sqrt{2}a=22"

"2a+\sqrt{2}a=11"

"a=\frac{11}{2+\sqrt{2}}"

"a=\frac{11}{2}(2-\sqrt{2})"

Side of the square is "\frac{11}{2}(2-\sqrt{2})."